OpenStudy (anonymous):

if cosx=3/5 then cos2x=? a)-7/25 b)6/10 c)4/5 d)12/25

OpenStudy (anonymous):

cos2x=cos^2x-sin^2x=cos^2x-1+cos^2x=2cos^2x-1

OpenStudy (anonymous):

cos2x=2cos^2 x-1=2(3/5)^2-1=-7/25

OpenStudy (anonymous):

thank you!